# FacAID: A Transformer Model for Neuro-Symbolic Facade Reconstruction

### v1

@@ Abstract.
We introduce a neuro-symbolic transformer-based model that converts flat, segmented facade structures into procedural definitions using a custom-designed split grammar. To facilitate this, we first develop a semi-complex split grammar tailored for architectural facades and then generate a dataset comprising of facades alongside their corresponding procedural representations. This dataset is used to train our transformer model to convert segmented, flat facades into the procedural language of our grammar. During inference, the model applies this learned transformation to new facade segmentations, providing a procedural representation that users can adjust to generate varied facade designs. This method not only automates the conversion of static facade images into dynamic, editable procedural formats but also enhances the design flexibility, allowing for easy modifications and variations by architects and designers. Our approach sets a new standard in facade design by combining the precision of procedural generation with the adaptability of neuro-symbolic learning.
@@
ccs: Computing methodologies Shape modelingccs: Computing methodologies Neural networks
@@ 1. Introduction
The field of architectural modeling in computer graphics has greatly advanced with the adoption of procedural methods, enabling the efficient generation of vast and complex urban structures and environments (Wonka et al., 2003; Müller et al., 2006). However, the creation of grammar rules and their parameters is time-consuming and requires specific knowledge and skills, posing a barrier for non-experts. In response, researchers in computer graphics introduced inverse procedural modeling to extract procedural rules from existing models or images, simplifying the process and making it more accessible (Musialski and Wimmer, 2013; Wu et al., 2014; St’ava et al., 2010).
@@
Our model is trained on a dataset of segmented facade structures and their corresponding procedural definitions, aiding the model in understanding and replicating structural nuances while aligning with architectural principles. During inference, the model applies learned transformations to new segmented images, converting them into procedural formats. This allows users to generate variations of designs, adjust elements, or create new facades within an established procedural framework, enhancing design flexibility and creativity. Figure 1 shows a urban scene designed by a digital artist using facades generated with our system.
The key contributions of our work include: (1) A neuro-symbolic transformer-based model that automates the conversion of segmented facades into editable procedural definitions. (2) A custom-designed split grammar specifically for architectural facades. (3) A novel dataset of facade structure and their procedural definitions, serving as a training and validation foundation. (4) Demonstrations of the model’s practical application in generating varied facade designs, showcasing its utility.
@@ 2. Related Work
Our work builds on grammar-based procedural methods, introduced for modeling of geometry in botany (Prusinkiewicz and Lindenmayer, 1990) and design (Stiny, 1975). For modeling of architecture, shape grammars (Stiny, 1975) based on splitting rules turned out to be well suitable (Wonka et al., 2003; Müller et al., 2006; Whiting et al., 2009). Such rules decompose geometric shapes into hierarchical structures, making it easier to manage and manipulate complex designs. The primary objective of our work is to automatically infer proper production rules from a grammar given an input segmented facade, which allows for user-friendly editing and generation of new variations (Lipp et al., 2008; Bao et al., 2013; Ilčík et al., 2015).
@@
Inverse procedural modeling aims to derive shape grammars from given segmented geometry or imagery. Early work introduced grammars that split building but each rule was crafted and applied manually (Bekins and Aliaga, 2005; Aliaga et al., 2007). Subsequent research has continued along this path, proposing various methods to extract procedural descriptions from facades (Müller et al., 2007; Becker, 2009; Teboul et al., 2011, 2013; Demir and Aliaga, 2018). Notably, the work on inverse procedural modeling of vector-art provided a formal treatment of the problem in graphics (St’ava et al., 2010).
A significant challenge in inverse procedural modeling is dealing with noisy or unsegmented inputs, where lower-level shape understanding and symmetry detection become crucial (Müller et al., 2007; Bokeloh et al., 2010). Once shape grammars are learned from typical input facades, they can be used as priors to guide further reconstruction efforts, a line of work explored by several research groups (Mathias et al., 2011; Ripperda and Brenner, 2009; Vanegas et al., 2010; Toshev et al., 2010; Riemenschneider et al., 2012; Teboul et al., 2013).
There have been multiple approaches to extract grammars from simple facades (Weissenberg et al., 2013), also with varying alternative subdivision rules (Zhang et al., 2013). Additionally, the concept of Bayesian models merging to combine deterministic grammars into a stochastic one has been explored, but it necessitates a robust initial grammar (Talton et al., 2012; Martinovic and Gool, 2013), or a predefined grid-layout structures (Fan et al., 2014). Learning of grammars based on the smallest-grammar paradigm (Charikar et al., 2005) using approximate dynamic programming (Wu et al., 2013) has also been proposed. In contrast, our approach does rely on pre-defined rules based on a split grammar (Wonka et al., 2003), but it uses neuro-symbolic transformer model to convert all segmented facade images into procedural representations robustly, streamlining the process of rule applications.
@@
Generative models have advanced significantly in recent years, particularly in modeling 2D images (Karras et al., 2018) and 3D shapes (Nash et al., 2020). These models often use latent space projections to generate new data (Richardson et al., 2020). In the context of building facade images, generative models create realistic urban images, which, although detailed, are often limited in resolution and lack high-level structure or parameterization (Bachl and Ferreira, 2019; Sun et al., 2022). Our work addresses these shortcoming by focusing on generating procedural graphs, allowing for enhanced flexibility and customization in facade design.
@@ 3. Problem Statement and Overview
During the inference, segmented facade abstractions are passed as input to the transformer model. It then predicts a procedure that correlates with the underlying structure of the input facade. After this step, a hierarchical procedural representation of all the facade elements is generated, but the sizing parameters are not yet set at this stage. As the created procedure is differentiable wrt the sizing parameters, a second module—the optimizer—minimizes the MSE between the target and the procedural result on pixel level basis.
@@ 4. Procedural Data Generation
To our best knowledge, there are no publicly available datasets that contain facades annotated with their procedural definitions of any kind. To create a dataset with facades and their procedural definitions, we developed a split grammar system, drawing on facade modeling techniques from previous work (Wonka et al., 2003). Our grammar includes around 30 production rules that range from simple symbol replacements to complex pattern formations and involves 16 terminal and 12 non-terminal symbols. Please refer to Figure 3 and supplemental material for visual depiction.
A random derivation of the grammar’s rule-set will still generate a non-plausible facade in the vast majority of cases. To address this, we implemented a generator that is able to generate a varied distribution of coherent facades based on architectural rules and heuristics. The generation process begins with a selection from style-specific subsets of arguments to maintain consistency. We define the chosen style-specific subset as $S$. The generator is initialized with a starting non-terminal $n_{0}$ and employs a heuristic module $H_{p}(S)$ to pick the next production to apply while maintaining architectural suitability. Another module, $H_{a}(S)$, adjusts the arguments of the selected production rule for structural soundness and coherence with the rest of the facade. This sequence of applying rules and adjusting arguments continues until all non-terminals are transformed into terminals, leaving a complete facade and a derivation tree that documents the creation process, and which can be executed in order to regenerate the facade.
It is those trees of productions that form the language of the procedures we focus on generating. Because of the nature of their structure, they are similar to other node-based procedural systems like shader graphs, but instead of performing mathematical calculations, each node partitions the space provided as the input to the node.
@@ 5.1. Transformer-Based Inverse Procedural Learning
In order to feed the data through our model, the inputs and outputs need to be converted into sequences of tokens, respectively $S_{I}$ and $S_{O}$. The input sequence represents a facade segmentation comprised of non-overlapping rectangles of different classes. Each rectangle $r_{i}$ is represented by 5 tokens $(t_{i},o^{x}_{i},o^{y}_{i},w_{i},h_{i})$ where $t_{i}$ indicates the grammar terminal symbol which the rectangle represents, $o^{x}_{i}$ and $o^{y}_{i}$ are the coordinates of its bottom-left corner and $w_{i}$ and $h_{i}$ are respectively its width and height. All of the rectangles $r_{i}$ are then sorted based on the position of their bottom-left corners (first by the $Y$ coordinate then $X$), in order to decrease the variability of the input sequences that represent the same facades, and concatenated to form the sequence $S_{I}$. Formulating the output procedures as sequences $S_{O}$ posses a greater challenge. Since the procedures take the form of rooted trees we employ the breadth-first traversal method (starting from the root) to produce a sequence of nodes $n_{i}$. After that we can resume with a similar approach as with the inputs - each node $n_{i}$ is represented by its class $c_{i}$ and the set of its structural parameters $(p^{1}_{i},...,p^{j}_{i})$ which can be either integer of floating point values. The nodes $n_{i}$ are then concatenated with separator tokens in-between them. When undergoing the tokenization process both categorical and integer parameters are translated exactly into their token counterparts whereas continuous values, like positions and sizes, are discretized between their edge values with a given resolution. After that the model can be trained on all the generated input-output sequence pairs of facade segmentations and their procedural representations.
@@
We employ the usual transformer approach and feed the model with not only the sequences of tokens but also the information about their positions in the sequence. Each token is tagged with its global position in either the input or output sequence. Additionally, since our rectangle sequences are represented by groups of 5 tokens we assign each of them a local index from 0 to 4 signifying their location in the sequence of 5 (Leimer et al., 2022). We extend this approach to the output sequences and assign the output tokens a local index in groups that encompass a single production and its arguments. In this case the groups are of dynamic length.
@@
Figure 4. Sizing optimizer: The input to sizing optimization is the reconstructed procedure (structural) with default sizing information. As our derived procedures are differentiable wrt the sizing parameters, the terminal sizes are computed by optimizing the square error.
@@ 5.2. Neuro-Symbolic Procedure Inference
The neuro-symbolic generation (inference) is performed by providing a segmented facade image, which can be done using interactive methods (Musialski et al., 2012) or by using automatic facade segmentation (Tyleček and Šára, 2013) or building labeling methods (Selvaraju et al., 2021). In the following we consider the input facade segmentation is given and represented as a flat irregular 2D grid of boxes with their classes (e.g., Window, Balcony, etc.) and extents.
@@
Figure 6. Optimized facades classification error distribution: Showcases how closely the optimized result of our pipeline matches with the input segmentation. The error measures what percentage of the resulting segmentation does not perfectly overlap the input.
Table 1. The average negative log likelihood loss of our model based on a test set of 10k facades. Both rows show the test loss obtained using models trained on different amounts of data (50k, 100k and 200k facades respectively) with and without Invalid State Nullification (refer to Section 5.2).
@@ 5.3. Differentiable Procedure Sizing Optimization
The model inference process leaves us with a structurally sound representation of the input facade but all of its sizing parameters are set to default values. In order to fully reproduce the input facade the procedure needs to go through the optimization module which finds the proper values of the sizing parameters. This module heavily leverages the fact that we have implemented all of our production rules in a way that makes them differentiable with respect to their sizing parameters. This means that when the procedure is executed all positions and sizes of resulting rectangles can be expressed as differentiable functions of the parameters of the procedure. We use this property to perform a gradient based optimization.
In order to perform the optimization we define a loss function that enables the model to sequentially approach the input facade segmentation. We treat the input segmentation as an image and then calculate the mean square error between it and the result of executing our procedure (Figure 4). The engine result also needs to be rasterized since it’s a a sequence of rectangles. To keep the function differentiable we rasterize all of the rectangles in a soft manner—using a combination of sigmoid functions positioned at the edges of the rectangles. This allows us to calculate the gradient of the mean square error with respect to all of the sizing parameters of the procedure.
@@
(b) Resolution of the tokenization. Our model is mostly agnostic wrt to that.
@@ 5.4. Implementation and Training
We have implemented our system using Python and the PyTorch library. We trained our model using datasets of varying sizes (50k, 100k, 200k) on a NVIDIA L4 24GB and NVIDIA A100 40GB graphics cards. The training time for 50k samples was 13h on the L4, and 2.8 times faster on the A100 (5h, 10h, 20h respectively for each training data size). For most of our subsequent quantitative results we used the 50k dataset, as it has the same convergence behavior as the bigger sets (cf. Fig 7(a)), however, the qualitative results presented below were generated using the 200k samples dataset.
@@ 6. Quantitative Evaluation
6.1. Procedure Reconstruction Frequency
@@
6.2. Procedure Structural Difference: Tree Edit Distance
The structural reconstructions are also measured against the ground truth based on the number of edits required in order to arrive at a precise reconstruction of the ground truth derivation tree. We use tree edit distance as our metric of the edits assuming that a single edit can amount to the following changes in the tree: (1) removing a node, (2) adding a node, (3) relabeling a node (changing its production type). Figure 5(b) shows that the majority of our structural reconstructions does not require any edits, while a vast majority requires no more than 5 editing steps. This, in most cases, means that less than 10% of the total tree needs to be modified to arrive at the ground truth structure.
6.3. Reconstruction Classification Error
We perform the quantitative measure of the optimized module on 1k, of the generated 10k, facade structures. The structures are optimized against a target segmentation from the ground truth. The metric used to compare the result with the input facade calculated what percentage of the segmentation does not match the ground truth exactly. In most cases the optimizer is able to adjust the parameters of the structural representations almost exactly, while a vast majority of the facades has an at least 80% match with the ground truth (cf. Fig. 6).

### v2

# FaçAID: A Transformer Model for Neuro-Symbolic Facade Reconstruction

@@ Abstract.
We introduce a neuro-symbolic transformer-based model that converts flat, segmented facade structures into procedural definitions using a custom-designed split grammar. To facilitate this, we first develop a semi-complex split grammar tailored for architectural facades and then generate a dataset comprising of facades alongside their corresponding procedural representations. This dataset is used to train our transformer model to convert segmented, flat facades into the procedural language of our grammar. During inference, the model applies this learned transformation to new facade segmentations, providing a procedural representation that users can adjust to generate varied facade designs. This method not only automates the conversion of static facade images into dynamic, editable procedural formats but also enhances the design flexibility, allowing for easy modifications.
@@
submissionid: 817journalyear: 2024copyright: rightsretainedconference: SIGGRAPH Asia 2024 Conference Papers; December 3–6, 2024; Tokyo, Japanbooktitle: SIGGRAPH Asia 2024 Conference Papers (SA Conference Papers ’24), December 3–6, 2024, Tokyo, Japandoi: 10.1145/3680528.3687657isbn: 979-8-4007-1131-2/24/12ccs: Computing methodologies Shape modelingccs: Computing methodologies Neural networks
@@ 1. Introduction
The field of architectural modeling in computer graphics has greatly advanced with the adoption of procedural methods, enabling the efficient generation of vast and complex urban structures and environments (Wonka et al., 2003; Müller et al., 2006). However, the creation of grammar rules and their parameters is time-consuming and requires specific knowledge, posing a barrier for non-experts. In response, researchers in computer graphics introduced inverse procedural modeling to extract procedural rules from existing models or images, simplifying the process and making it more accessible (Musialski and Wimmer, 2013; Wu et al., 2014; St’ava et al., 2010).
@@
Our model is trained on a dataset of segmented facade structures and their corresponding procedural definitions, aiding the model in understanding and replicating structural nuances while aligning with architectural principles. During inference, the model applies learned transformations to new segmented images, converting them into procedural formats. This allows users to generate variations of designs, adjust elements, or create new facades within an established procedural framework, enhancing design flexibility and creativity. Figure 1 shows an urban scene designed by a digital artist using facades generated with our system.
The key contributions of our work include: (1) A neuro-symbolic transformer-based model that automates the conversion of segmented facades into editable procedural definitions. (2) A custom-designed split grammar specifically for architectural facades. (3) A novel dataset of facade structures and their procedural definitions, serving as a training and validation foundation. (4) Demonstrations of the model’s practical application in generating varied facade designs, showcasing its utility.
@@ 2. Related Work
Our work builds on grammar-based procedural methods, introduced for modeling of geometry in botany (Prusinkiewicz and Lindenmayer, 1990) and design (Stiny, 1975). For modeling of architecture, shape grammars (Stiny, 1975) based on splitting rules turned out to be well suitable (Wonka et al., 2003; Müller et al., 2006; Whiting et al., 2009). Such rules decompose geometric shapes into hierarchical structures, making it easier to manage complex designs. The primary objective of our work is to automatically infer production rules from a grammar given an input segmented facade, which allows for user-friendly editing and generation of new variations (Lipp et al., 2008; Bao et al., 2013; Ilčík et al., 2015).
@@
Inverse procedural modeling aims to derive shape grammars from given segmented geometry or imagery. Early work introduced grammars that split buildings but each rule was crafted and applied manually (Bekins and Aliaga, 2005; Aliaga et al., 2007). Subsequent research has continued along this path, proposing various methods to extract procedural descriptions from facades (Müller et al., 2007; Becker, 2009; Teboul et al., 2011, 2013; Demir and Aliaga, 2018). Notably, the work on inverse procedural modeling of vector-art provided a formal treatment of the problem in graphics (St’ava et al., 2010). A significant challenge in inverse procedural modeling is dealing with noisy or unsegmented inputs, where lower-level shape understanding and symmetry detection become crucial (Müller et al., 2007; Bokeloh et al., 2010). Once shape grammars are learned from typical input facades, they can be used as priors to guide further reconstruction efforts (Mathias et al., 2011; Ripperda and Brenner, 2009; Vanegas et al., 2010; Toshev et al., 2010; Riemenschneider et al., 2012b; Teboul et al., 2013).
There have been multiple approaches to extract grammars from simple facades (Weissenberg et al., 2013), also with varying alternative subdivision rules (Zhang et al., 2013). Additionally, the concept of Bayesian models merging to combine deterministic grammars into a stochastic one has been explored, but it necessitates a robust initial grammar (Talton et al., 2012; Martinovic and Gool, 2013), or predefined grid-layout structures (Fan et al., 2014). Learning of grammars based on the smallest-grammar paradigm (Charikar et al., 2005) using approximate dynamic programming (Wu et al., 2013) has also been proposed. In contrast, our approach relies on pre-defined rules based on a split grammar (Wonka et al., 2003), but it uses neuro-symbolic transformer model to convert all segmented facade images into procedural representations robustly, streamlining the process of rule applications.
@@
Generative models have advanced significantly in recent years, particularly in modeling 2D images (Karras et al., 2018) and 3D shapes (Nash et al., 2020). These models often use latent space projections to generate new data (Richardson et al., 2020). In the context of building facade images, generative models create realistic urban images, which, although detailed, are often limited in resolution and lack high-level structure or parameterization (Bachl and Ferreira, 2019; Sun et al., 2022). Our work addresses these shortcomings by focusing on generating procedural graphs, allowing for enhanced flexibility and customization in facade design.
@@ 3. Problem Statement and Overview
During inference, segmented facade abstractions are passed as input to the transformer model, which predicts a procedure correlating with the underlying structure of the input facade. After this step, a hierarchical procedural representation of all facade elements is generated, but the sizing parameters are not yet set. As the created procedure is differentiable w.r.t. the sizing parameters, a second module—–the optimizer—–minimizes the MSE between the target and the procedural result on a pixel level basis.
@@ 4. Procedural Data Generation
To the best of our knowledge, there are no publicly available datasets that contain facades annotated with their procedural definitions of any kind. To create a dataset with facades and their procedural definitions, we developed a split grammar system, drawing on facade modeling techniques from previous work (Wonka et al., 2003). Please refer to Figure 3 and supplemental material for visual depiction. A random derivation of the grammar’s rule-set will still generate a non-plausible facade in the vast majority of cases. To address this, we implemented a generator that is able to produce a varied distribution of coherent facades based on architectural rules and heuristics.
The generation process begins with selecting style-specific subsets of architectural parameters. We define them as $S$. They serve as a guideline for keeping the generated design coherent. The generator is then initialized with a starting non-terminal. At each iteration the generator chooses from production rules applicable to the current non-terminal. It employs a heuristic module $H_{p}(S)$ to select the next proper production to apply. After that another heuristic module, $H_{a}(S)$, picks the arguments of the selected production rule. The two modules work together to maintain architectural suitability, structural soundness and coherence throughout the whole facade. This sequence of applying rules and adjusting arguments continues until all non-terminals are processed, leaving a complete facade and a derivation tree that documents the creation process, and which can be executed in order to regenerate the facade.
It is those trees of productions that form the language of the procedures we focus on generating. Because of their structure, they are similar to other node-based procedural systems like shader graphs, but instead of performing mathematical calculations, each node partitions the space provided as the input to the node.
The data and the procedural generator are available on the project’s website (facaid.github.io). Please refer to the code for any detail regarding the structure of the grammar as well as the heuristics.
@@ 5.1. Transformer-Based Inverse Procedural Learning
In order to feed the data through our model, the inputs and outputs need to be converted into sequences of tokens, respectively $S_{I}$ and $S_{O}$. The input sequence represents a facade segmentation comprised of non-overlapping rectangles of different classes. Each rectangle $r_{i}$ is represented by 5 tokens $(t_{i},o^{x}_{i},o^{y}_{i},w_{i},h_{i})$ where $t_{i}$ indicates the grammar terminal symbol which the rectangle represents, $o^{x}_{i}$ and $o^{y}_{i}$ are the coordinates of its bottom-left corner and $w_{i}$ and $h_{i}$ are respectively its width and height. All of the rectangles $r_{i}$ are then sorted based on the position of their bottom-left corners (first by the $Y$ coordinate then $X$), in order to decrease the variability of the input sequences that represent the same facades, and concatenated to form the sequence $S_{I}$.
Formulating the output procedures as sequences $S_{O}$ poses a greater challenge. Since the procedures take the form of rooted trees we employ the breadth-first traversal method (starting from the root) to produce a sequence of nodes $n_{i}$. After that we can resume with a similar approach as with the inputs—each node $n_{i}$ is represented by its class $c_{i}$ and the set of its structural parameters $(p^{1}_{i},...,p^{j}_{i})$ which can be integer or floating point values. The nodes $n_{i}$ are then concatenated with separator tokens in-between them.
When undergoing the tokenization process both categorical and integer parameters are translated into their token counterparts whereas continuous values, like positions and sizes, are discretized between their edge values with a given resolution. After that the model can be trained on all the generated input-output sequence pairs of facade segmentations and their procedural representations.
@@
We employ the usual transformer approach and feed the model with not only the sequences of tokens but also the information about their positions in the sequence. Each token is tagged with its global position in the sequence. Additionally, since our rectangle sequences are represented by groups of 5 tokens we assign each of them a local index from 0 to 4 signifying their location in the sequence of 5 (Leimer et al., 2022). We extend this approach to the output sequences and assign the output tokens a local index in groups that encompass a single production and its arguments. In this case the groups are of dynamic length.
@@
(a) Starting facade
(b) Target facade
(c) Square error
(d) Optimized facade
Figure 4. Sizing optimizer: The input to sizing optimization is the reconstructed procedure (structural) with default sizing information. As our derived procedures are differentiable w.r.t. the sizing parameters, the terminal sizes are computed by optimizing the square error.
@@ 5.2. Neuro-Symbolic Procedure Inference
The neuro-symbolic generation (inference) is performed by providing a segmented facade image, which can be done using interactive methods (Musialski et al., 2012) or by using automatic facade segmentation (Tyleček and Šára, 2013) or building labeling methods (Selvaraju et al., 2021). We consider the input facade segmentation is given and represented as a flat irregular 2D grid of boxes with their classes (e.g., Window, Balcony, etc.) and extents.
@@
(a) Optimized facades classification error distribution. The error compares the input and output segmentations (pixelwise).
(b) Average tree edit distance to the ground truth when performing inference using inputs with a given level of noise.
Figure 6. Facade reconstruction error
Table 1. The average NLL loss of our model on a test set of 10k facades. Both rows show the test loss obtained using models trained on different amounts of data (50k, 100k and 200k respectively) with and without Invalid State Nullification (refer to Section 5.2).
@@ 5.3. Differentiable Procedure Sizing Optimization
The model inference process leaves us with a structurally sound representation of the input facade but all of its sizing parameters are set to default values. In order to fully reproduce the input facade the procedure needs to go through the optimization module which finds the proper values of the sizing parameters. This module heavily leverages the fact that we have implemented all of our production rules in a way that makes them differentiable w.r.t. their sizing parameters. This means that when the procedure is executed all positions and sizes of resulting rectangles can be expressed as differentiable functions of the parameters of the procedure. We use this property to perform a gradient based optimization.
In order to perform the optimization we define a loss function that enables the model to sequentially approach the input facade segmentation. We treat the input segmentation as an image and then calculate the mean square error between it and the result of executing our procedure (Figure 4). The engine result also needs to be rasterized since it’s a sequence of rectangles. To keep the function differentiable we rasterize all of the rectangles in a soft manner—using a combination of sigmoid functions at the edges of the rectangles. This allows us to calculate the gradient of the mean square error w.r.t. all of the sizing parameters of the procedure.
@@
(b) Resolution of the tokenization. Our model is mostly agnostic w.r.t. to that.
@@ 5.4. Implementation and Training
We have implemented our system using Python and the PyTorch library. We trained our model using datasets of varying sizes (50k, 100k, 200k) on a NVIDIA L4 24GB and the NVIDIA A100 40GB graphics cards. The training time for 50k samples was 13h on the L4, and 2.8 times faster on the A100 (5h, 10h, 20h respectively for each training data size). For most of our subsequent quantitative results we used the 50k dataset, as it has the same convergence behavior as the bigger sets (cf. Fig 7(a)), however, the qualitative results presented below were generated using the 200k samples dataset. Additional material is available on facaid.github.io.
@@ 6. Quantitative Evaluation
Procedure Reconstruction Frequency
@@
Procedure Structural Difference: Tree Edit Distance
The structural reconstructions are also measured against the ground truth based on the number of edits required to arrive at a precise reconstruction of the ground truth derivation tree. We use tree edit distance as our metric of the edits assuming that a single edit can amount to the following changes in the tree: (1) removing a node, (2) adding a node, (3) relabeling a node (changing its production type). Figure 5(b) shows that the majority of our structural reconstructions does not require any edits, while a vast majority requires no more than 5 editing steps. This, in most cases, means that less than 10% of the total tree needs to be modified to arrive at the ground truth structure.
Reconstruction Classification Error
We perform the quantitative measure of the optimized module on 1k, of the generated 10k, facade structures. The structures are optimized against a target segmentation from the ground truth. The metric used to compare the result with the input facade calculated what percentage of the segmentation does not match the ground truth exactly. In most cases the optimizer is able to adjust the parameters of the structural representations almost exactly, while a vast majority of the facades has an at least 80% match with the ground truth (cf. Fig. 6(a)).
Robustness to Noise in the Input
We have tested the robustness of our model when provided with noisy segmentations. We generate a test set of 500 procedural facades, never seen before by the model. After that we apply different levels of noise to all elements of each segmentation. The level of noise is calculated based on the percentage of pixels in the new segmentation which differ from the original one. We check the quality of the structural reconstruction of our model by comparing the inference results generated from the noisy inputs with the ground truth procedures using tree edit distance. Figure 6(b) showcases the average tree edit distance for a given noise level. The model is able to reconstruct segmentations with up to 10% noise with a good level of accuracy—less than 6 edits to arrive at the ground truth.
7. Qualitative Results
7.1. Comparison to Previous Methods
We compare the results of our method to the work of Riemenschneider et al. (2012a) and Teboul et al. (2013) by using the same target facades and segmenting them by hand in a similar way as as shown in their papers. Figure 8 showcases the respective results. Our method demonstrates the capability to reproduce procedural representations which are comparable to their results. Note that the results are highly dependent on the input segmentations, the creation of which can be ambiguous: if we include the balconies explicitly in the segmentation they are recognized by our algorithm; if not, they are missing in our result, producing the same output as their method (Figure 8(b)). Note that their results were taken directly from the papers because of the lack of available implementations.
Input
Refer to caption     Refer to caption
  Theirs
  Ours
(a) (Riemenschneider et al., 2012a)
Refer to captionRefer to captionRefer to caption
Refer to captionRefer to caption
(b) (Teboul et al., 2013)
Figure 8. Comparison to the work of Riemenschneider et al. (2012a) and Teboul et al. (2013). The first column shows the original facade image and its segmentation(s). The second column displays the results of their methods. The last column shows the segmented facades produced by our method.
7.2. Qualitative Evaluation
The qualitative results were generated using manually segmented and annotated facades. The segmentations have been crafted such that they match the terminals of our grammar and take around 5 minutes each to construct. For that task we have implemented a rudimentary segmentation interface. In practice, our goal would be to either use automatic segmentation tools, based on machine learning methods (Selvaraju et al., 2021) or on interactive segmentation tools (Musialski et al., 2012). However, generally, we consider that as given input which is not part of and essentially does not influence the contribution of this paper.